The square root of (6-x) + The square root of (5x + 6) = 6
Answer by stanbon(75887)

square root of (6-x) + The square root of (5x + 6) = 6
Square both sides to get:
(6-x)+2sqrt[(6-x)(5x+6)]+5x+6 = 36
2sqrt[......] =24-4x
sqrt[....]=12-2x
Square both sides again to get:
-(5x+6)(x-6)=[-2(x-6)]^2
-[5x^2-24x-36] = 4[x^2-12x+36]
-5x^2+24x+36 = 4x^2-48x+144
-9x^2+72x-108=0
x^2-8x+12=0
(x-6)(x-2)=0
x=6 or x=2
Checking the answers for extraneous roots:
If x=6 the original equation is true
If x=2 the original equation is true
=================
cheers,
Stan H.